# The polynomial set associated with a fixed number of matrix-matrix multiplications

### v1

@@ 5.3 Six multiplications
To our knowledge, state-of-the-art for $m=6$ multiplications is $d=24$, again given in [25], with $s=3$. In our terminology, that corresponds to the reduction $a_{23}=a_{33}=a_{34}=a_{67}=0$. Similar to the situation for $m=5$, this is not the highest admissable degree. From Figure 1 we see that the highest admissible degree is $d=32$. With $r=3$, we can use the reduction (62).
@@
Unfortunately, the application of the package HomotopyContinuation.jl was not successful for this case. We have $33$ equations, and $m^{2}-r=33$ unknowns. The creation of initial vectors for the homotopy continuation seems too computationally demanding, possibly due to the high total degree of the polynomial system. Instead we used the package GraphMatFun.jl [14] to create a locally convergent iterative solution method. The system is highly ill-conditioned and a standard Newton approach was not successful. Instead, we employ an iteratively regularized Newton’s method. Following the approach of [6], we compute a Tikhonov-Newton step by applying Newton’s method to a Tikhonov-regularized system. The Tikhonov-Newton step can be computed using the singular value decomposition of the Jacobian matrix, which is directly available from the graph representation in GraphMatFun.jl. The best results were obtained by using Armijo steplength damping and selecting between a Newton and a Tikhonov-Newton step. For the sake of reproducibility, the starting vectors are given explicitly in the software available in the github repository. In order determine conclusively that a solution is found, the solution was post-processed with high precision arithmetic (BigFloat in Julia) such that the 50 first decimals of the coefficients are seemingly accurate.
@@
The above approach with the structure (62) was successful in finding a solution vector for the problem corresponding to the Taylor expansion of the matrix exponential, using complex arithmetic. Unfortunately, we were not able to find a real coefficient vector. The maximum absolute number in the coefficient vector is $3.7\cdot 10^{5}$, and the Jacobian evaluated at the solution is $5.8\cdot 10^{11}$.
@@
From an application viewpoint, real coefficients are advantageous, e.g., since the evaluation of $p(A)$ can be done in real arithmetic if $A$ is real. In order to find a real evaluation scheme, we investigated instead the following structure:
$[A|B]=\left[\begin{array}[]{ccccccc|ccccccc}0&1&&&&&&0&1&&&&\\ 0&\times&1&&&&&0&\times&1&&&\\ 0&\times&\times&1&&&&0&\times&1&0&&\\ 0&\times&\times&\times&1&&&0&\times&\times&\times&1&\\ 0&\times&\times&\times&\times&1&&0&\times&\times&\times&1&0\\ 0&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&1&0\\ \end{array}\right].$
@@
This leads to a polynomial of degree $d=30$. Using the locally convergent iterative method described above, we successfully computed a real solution vector. The maximum absolute number in the coefficient vector is $5.0\cdot 10^{4}$, and the condition number of the Jacobian at the solution is $2.10\cdot 10^{10}$.
@@
The highest degree polynomial set can indeed be different for real and complex case, and the simulations suggest the following.
Conjecture 5.2 (Six multiplications).
@@ 5.4 Seven multiplications
For seven multiplications, we believe state-of-the-art is $d=30$, [25, Table 3] with $s=4$, which in our context corresponds to the reduction $a_{23}=a_{33}=a_{34}=a_{43}=a_{44}=a_{45}=0$ and $a_{78}=0$. In Figure 1, we see that the highest admissible degree seems to be $d=42$ with $r=5$. Using the same technique as for $m=6$ we were able to find a solution to the system for the Taylor expansion of the exponential, if we input scaling $\alpha=16$ and when we used complex arithmetic for the highest admissible degree. We used the following structure:
$[A|B]=\left[\begin{array}[]{cccccccc|ccccccccc}0&1&&&&&&&0&1&&&\\ 0&\times&1&&&&&&0&\times&1&&\\ 0&\times&0&1&&&&&0&\times&1&0&\\ 0&\times&\times&\times&1&&&&0&\times&\times&1&0&\\ 0&\times&0&\times&\times&1&&&0&\times&\times&\times&1&0&\\ 0&\times&\times&\times&\times&\times&1&&0&\times&\times&\times&\times&1&0&\\ 0&\times&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&\times&1&0&\\ \end{array}\right].$
The fixed elements in $a_{3,3}=a_{5,3}=0$ were selected empirically to keep the condition number low. The maximum absolute number in the coefficient vector is $2.8\cdot 10^{6}$, and the condition number of the Jacobian at the solution is $2.9\cdot 10^{13}$. From this we conclude the following in a complex sense.
Conjecture 5.3 (Seven multiplications).
@@ 6 Conclusions
This work has a characterization of the set $\Pi_{2^{m}}^{*}$ as its focus, with particular attention given to minimality and computing the maximum degree polynomial subset of $\Pi_{2^{m}}^{*}$ in the sense of (64). The minimality question is from our perspective rather well understood in this paper. The determination of the maximum degree polynomial subset can be further investigated. We only used described $m\leq 7$ with computational reasoning. Both the theoretical description of the general case (e.g. using further tools from algebraic geometry [32]), as well as a more general computational approach could be of interest and useful in practice.
@@
Based on our simulations, one major component is missing before this can be directly used in matrix function evaluation software: Understanding of rounding errors. Evaluation of high-degree polynomials are in general prone to rounding errors, unless special representations such as a Chebyshev basis is used. In this case it seems even more intricate. For example, the coefficients we computed, were computed to full double precision using high precision arithmetic. However, the fact that the system has a rather large condition number, is an indication that in general this evaluation is sensitive with respect to these coefficients. Heuristics similar to [15] might be applicable in a general setting. Further work is needed to determine which evaluation schemes, in the continuuum of $(A,B,c)$ lead to better numerical stability, e.g., in the sense that the condition number is smaller.
@@
The Paterson-Stockmeyer method has proven beneficial not only for evaluating matrix polynomials, but similar computational challenges can arise in other contexts, such as when the input $X$ is a polynomial. In these scenarios, multiplying two quantities is significantly more computationally demanding than forming linear combinations. The construction in the Paterson-Stockmeyer approach resembles the baby-step giant-step (BSGS) technique introduced in [30], which has found various applications and combined with Paterson-Stockmeyer approach in public key and privacy-preserving cryptography [12]. Moreover, both the Paterson-Stockmeyer method and BSGS serve as valuable tools in high-precision arithmetic [16, 31]. Open research questions include how the approach presented in this paper, or methods for $\Pi_{2^{m}}^{*}$ in general, can be applied in these contexts.
@@
The fixed-cost computation approach presented in this paper can be complemented by insights from research on composite polynomials or deep polynomials. See [20] for composite polynomials. Rational approximations corresponding to this concept, such as those in [10], illustrate how successive compositions, e.g., $p(f(g(x)))$, can achieve rapid convergence in terms of both the number of compositions and the parameters involved. Similar findings are noted in [33], motivated by the link between this approach and universal approximation in deep learning. The composite polynomial method can fit within the framework of this paper by zeroing certain elements in matrices $A$ and $B$. Nevertheless, there is a significant distinction in research objectives: our objective is to minimize the number of matrix-matrix multiplications, while [33] focuses on reducing the number of parameters. Although some results, such as the approximation of the $p$th root [10], may be directly applicable, further research is needed to fully explore the differences resulting from these different objectives.
@@ Acknowledgements
The authors are grateful for the discussion and comments from Katlén Kohn, in particular concerning Section 2.2.

### v3

@@ 5.3 Six multiplications
To our knowledge, the state of the art for $m=6$ multiplications is $d=24$, again given in [25] with $s=3$. In our terminology, that corresponds to the reduction $a_{23}=a_{33}=a_{34}=a_{67}=0$. Similar to the situation for $m=5$, this is not the highest admissible degree. From Figure 1 we see that the highest admissible degree is $d=32$. With $r=3$, we can use the reduction (69).
@@
Unfortunately, the application of the package HomotopyContinuation.jl was not successful for this case. We have $33$ equations, and $m^{2}-r=33$ unknowns. The creation of initial vectors for the homotopy continuation seems too computationally demanding, likely related to the high total degree of the polynomial system. Instead we used the package GraphMatFun.jl [14] to create a locally convergent iterative solution method. The system is highly ill-conditioned and a standard Newton approach was not successful. Instead, we employ an iteratively regularized Newton’s method. Following the approach in [6], we compute a Tikhonov–Newton step by applying Newton’s method to a Tikhonov-regularized system. The Tikhonov–Newton step can be computed using the singular value decomposition of the Jacobian matrix, which is directly available from the graph representation in GraphMatFun.jl. The best results were obtained by using Armijo step-length damping and selecting between a Newton and a Tikhonov–Newton step in a greedy fashion. For the sake of reproducibility, the starting vectors are given explicitly in the software available in the GitHub repository. In order to determine conclusively that a solution is found, the solution was post-processed with high-precision arithmetic (BigFloat in Julia) so that the first 50 decimals of the coefficients appear accurate. This was done for this particular simulation as well as for all subsequent simulations.
@@
The above approach with the structure (69) was successful in finding a solution vector for the problem corresponding to the Taylor expansion of the matrix exponential, using complex arithmetic. Unfortunately, we were not able to find a real coefficient vector.
@@
From an application viewpoint, real coefficients are advantageous, e.g., since the evaluation of $p(A)$ can be done in real arithmetic if $A$ is real. In order to find a real evaluation scheme, we investigated instead the structure
$A|B]=\left[\begin{array}[]{ccccccc|ccccccc}0&1&&&&&&0&1&&&&\\ 0&\times&1&&&&&0&\times&1&&&\\ 0&\times&\times&1&&&&0&\times&1&0&&\\ 0&\times&\times&\times&1&&&0&\times&\times&\times&1&\\ 0&\times&\times&\times&\times&1&&0&\times&\times&\times&1&0\\ 0&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&1&0\\ \end{array}\right].$
@@
This leads to a polynomial of degree $d=30$. Using the locally convergent iterative method described above, we successfully computed a real solution vector.
For the actual numbers, we refer to the software. The maximum absolute value in the coefficient vector for the complex case is $3.7\cdot 10^{5}$, and the Jacobian evaluated at the solution has a condition number of $5.8\cdot 10^{11}$. The maximum absolute value in the coefficient vector for the real case is $5.0\cdot 10^{4}$, and the condition number of the Jacobian at the solution is $2.10\cdot 10^{10}$.
@@
The highest degree polynomial set can indeed be different for the real and complex cases, and the simulations suggest the following.
Conjecture 12 (Six multiplications).
@@ 5.4 Seven multiplications
For seven multiplications, we believe that the state of the art is $d=30$ [25, Table 3] with $s=4$, which in our context corresponds to the reduction $a_{23}=a_{33}=a_{34}=a_{43}=a_{44}=a_{45}=0$ and $a_{78}=0$. In Figure 1 we see that the highest admissible degree appears to be $d=42$ with $r=5$. Using the same technique as for $m=6$, we were able to find a solution to the system for the Taylor expansion of the exponential. We used scaling $\alpha=16$, complex arithmetic, and the following structure:
(77) $A|B]=\left[\begin{array}[]{cccccccc|ccccccccc}0&1&&&&&&&0&1&&&\\ 0&\times&1&&&&&&0&\times&1&&\\ 0&\times&0&1&&&&&0&\times&1&0&\\ 0&\times&\times&\times&1&&&&0&\times&\times&1&0&\\ 0&\times&0&\times&\times&1&&&0&\times&\times&\times&1&0&\\ 0&\times&\times&\times&\times&\times&1&&0&\times&\times&\times&\times&1&0&\\ 0&\times&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&\times&1&0&\\ \end{array}\right].$
The fixed elements $a_{3,3}=a_{5,3}=0$ were selected empirically in an attempt to avoid very large condition numbers. The maximum absolute value in the coefficient vector is $2.8\cdot 10^{6}$, and the condition number of the Jacobian at the solution is $2.9\cdot 10^{13}$. From this we conclude the following in a complex sense.
Conjecture 13 (Seven multiplications).
@@
Remark 14 (Computational example).
Although the main contributions of this manuscript are theoretical in character, we also wish to point out the practical value of the simulations. For example, the output of the simulations concerning $m=7$ multiplications can be directly used to compute the matrix exponential, in a very competitive way. In fact, in theory (in the sense of number of floating point operations) the method is the fastest for large norm matrices, as far as we know. Although a complete computational study is beyond the scope of this paper, we provide (for a specific but random matrix) a comparison with the scaling-and-squaring algorithm [18], which is the most commonly method used in mathematical software for the matrix exponential. In the following we see that our approach is faster or more accurate, or both depending on viewpoint, than two valid parameter choices for the scaling-and-squaring method. What is marked as our method is the evaluation (77) with coefficients precomputed with the Tikhonov-Newton method.
julia> Random.seed!(0); setprecision(128); n=200; alpha=16;
julia> A=randn(Complex{BigFloat},n,n); A=20*A/norm(A);
julia> E1 = @btime exp16_deg42_bigfloat(A/alpha); # Our method
  58.648 s (719681983 allocations: 29.50 GiB)
julia> E2 = @btime exp_sas_6mult_1div(A); # standard version 1
  65.340 s (762205498 allocations: 31.24 GiB)
julia> E3 = @btime exp_sas_7mult_1div(A); # standard version 2
  80.014 s (860125595 allocations: 35.25 GiB)
julia> # Compute a reference solution with high precision
julia> expAref=mapreduce(i-> (A^i)/factorial(big(i)), +, 0:100);
julia> norm(expAref-E1); # Error for our method
9.25199599560488132729984839075891589943e-33
julia> norm(expAref-E2); # Error for standard implementation version 1
1.935440065066927257455720406200123063236e-30
julia> norm(expAref-E3); # Error for standard implementation version 2
1.14214602689126618911305625822908609328e-36
@@ 6 Conclusions
This work focuses on a characterization of the set $\Pi_{2^{m}}^{*}$, with particular attention given to minimality and to computing the maximum degree polynomial subset of $\Pi_{2^{m}}^{*}$ in the sense of (71). From our perspective, the minimality question is well understood in this paper. The determination of the maximum degree polynomial subset can be further investigated. We have only described the case $m\leq 7$ using computational reasoning. Both a theoretical description of the general case, e.g., using further tools from algebraic geometry [33], and a more general computational approach could be of interest and useful in practice.
@@
Based on our simulations, one major component is missing before this can be directly used in matrix function evaluation software: understanding the effect of rounding errors. Evaluating high-degree polynomials is, in general, prone to rounding errors—unless special representations such as a Chebyshev basis are used. In this case, the issue appears even more intricate. For example, by using high-precision arithmetic, the coefficients were computed such that we could guarantee correctness in full double precision. However, the fact that the system has a rather large condition number (at least for $m=7$) is an indication that this evaluation is sensitive with respect to these coefficients. Heuristics similar to [15] might be applicable in a general setting. Further work is needed to determine which evaluation schemes, in the continuum of $(A,B,c)$, lead to better numerical stability; a necessary condition for numerical stability is that the condition number is not too large.
@@
The Paterson–Stockmeyer method has proven beneficial not only for evaluating matrix polynomials; similar computational challenges arise in other contexts, such as when the input $X$ is a polynomial. In these scenarios, multiplying two quantities is significantly more computationally demanding than forming linear combinations. The construction in the Paterson–Stockmeyer approach resembles the baby-step giant-step (BSGS) technique introduced in [31], which has found various applications and has been combined with the Paterson–Stockmeyer approach in public key and privacy-preserving cryptography [12]. Moreover, both the Paterson–Stockmeyer method and BSGS serve as valuable tools in high-precision arithmetic [16, 32]. Open research questions include how the approach presented in this paper, or methods for $\Pi_{2^{m}}^{*}$ in general, can be applied in these contexts.
@@
The fixed-cost computation approach presented in this paper can be complemented by insights from research on composite polynomials or deep polynomials. See [20] for composite polynomials. Rational approximations corresponding to this concept, such as those in [10], illustrate how successive compositions, e.g., $p(f(g(x)))$, can achieve rapid convergence in terms of both the number of compositions and the parameters involved. Similar findings are noted in [34], motivated by the link between this approach and universal approximation in deep learning. The composite polynomial method can fit within the framework of this paper by zeroing certain elements in matrices $A$ and $B$. Nevertheless, there is a significant distinction in research objectives: our objective is to minimize the number of matrix-matrix multiplications, while [34] focuses on reducing the number of parameters. Although some results, such as the approximation of the $p$th root [10], may be directly applicable, further research is needed to fully explore the differences resulting from these objectives.
@@ Acknowledgements
The authors wish to express gratitude for the insightful discussions and feedback from Prof. Kathlén Kohn (KTH Royal Insititute of Technology), particularly regarding Section 2.2. This research was partially conducted during the first author’s sabbatical at EPFL / University of Geneva. The support of the hosts, Prof. Daniel Kressner and Prof. Bart Vandereycken, is greatly appreciated. The authors also greatly acknowledge the valuable comments from Prof. Massimiliano Fasi (University of Leeds) and Prof. Jorge Sastre (Polytechnic University of Valencia).
Appendix A Auxiliary material for the proof of Theorem 10
The proof is based on two technical lemmas. We note that several of these results hold for arbitrary choices of $(A,B,c)$, but some parts of the statements and some of the proofs are clearer when we assume the specific structure of the triplet $(A,B,c)$ given in (57), which is sufficient for the proof of Theorem 10. Note that for our example, $Q_{j}=A_{j}B_{j}$, $A_{j}=Q_{2}+Q_{j-1}$, and $B_{j}=Q_{2}+\cdots+Q_{j-1}$.
The first lemma states a formula for the degree when we apply a differentiation operator consisting of a linear combination of derivatives of the elements in $(A,B,c)$ but not of the last rows of $A$ and $B$.
Lemma 15.
Let $(A,B,c)$ be of the specific structure given in (57). Let $\mathcal{D}_{i}$ be an operator consisting of a linear combination of the elements of $A,B$ in rows $1,2,\ldots,i$. Assume that
$\deg(\mathcal{D}_{i}Q_{3})\leq\cdots\leq\deg(\mathcal{D}_{i}Q_{i+1})<\deg(\mathcal{D}_{i}Q_{i+2})$
and that $\deg(\mathcal{D}_{i}Q_{i+2})\geq 1$. Then, $\deg(\mathcal{D}_{i}Q_{i+2})<\cdots<\deg(\mathcal{D}_{i}Q_{m+2})$ and
(78) $\deg(\mathcal{D}_{i}p)=\deg(\mathcal{D}_{i}Q_{m+2})=2^{m-1}+\cdots+2^{i}+\deg(\mathcal{D}_{i}Q_{i+2}).$
Proof.
We prove the statement by induction, starting with row $i$. By applying the product rule and using $\mathcal{D}_{i}Q_{2}=0$, we obtain:
(79) $\mathcal{D}_{i}Q_{i+3}=(\mathcal{D}_{i}A_{i+3})B_{i+3}+A_{i+3}(\mathcal{D}_{i}B_{i+3})=\\ (\mathcal{D}_{i}Q_{i+2})C_{i+3}+A_{i+3}\left(\mathcal{D}_{i}Q_{3}+\cdots+\mathcal{D}_{i}Q_{i+1}\right)$
where $C_{i+3}=A_{i+3}+B_{i+3}$. Note that $\deg(C_{i+3})=\deg(A_{i+3})=2^{i}$. Therefore, on the right-hand side of equation (79), the degree of the first term is larger than that of the second term, given our assumption that $\deg(\mathcal{D}_{i}Q_{3})\leq\cdots\leq\deg(\mathcal{D}_{i}Q_{i+1})<\deg(\mathcal{D}_{i}Q_{i+2})$. Thus, we have:
(80) $\deg(\mathcal{D}_{i}Q_{i+3})=\deg(A_{i+3})+\deg(\mathcal{D}_{i}Q_{i+2})=2^{i}+\deg(\mathcal{D}_{i}Q_{i+2})>\deg(\mathcal{D}_{i}Q_{i+2}).$
This establishes the base step of our induction.
To proceed with the induction step, assume that the inequality holds for $j$ steps, i.e.,
$\deg(\mathcal{D}_{i}Q_{i+2})<\cdots<\deg(\mathcal{D}_{i}Q_{i+j+1}).$
Analogous to the derivation in equation (79), we have:
$\mathcal{D}_{i}Q_{i+j+2}=(\mathcal{D}_{i}Q_{i+j+1})C_{i+j+2}+A_{i+j+2}\left(\mathcal{D}_{i}Q_{3}+\cdots+\mathcal{D}_{i}Q_{i+j}\right).$
Since $C_{i+j+2}$ and $A_{i+j+2}$ have the same degree, and by our induction assumption, the first term has a higher degree, we conclude that:
(81) $\deg(\mathcal{D}_{i}Q_{i+j+2})=\deg(C_{i+j+2})+\deg(\mathcal{D}_{i}Q_{i+j+1})=2^{i+j-1}+\deg(\mathcal{D}_{i}Q_{i+j+1}).$
This completes the proof of the increasing degree progression. The relation (78) follows by applying equation (81) for $j=m-i,\ldots,2$ and using (80) once. ∎
The previous lemma (Lemma 15) gives us the degree of the partial derivative of the output, given the degree of the partial derivative of $Q_{i+2}$. It remains to determine the derivatives of $Q_{i+2}$. We select the differentiation operator in several ways for the purpose of later combining them to form distinct degrees of the Jacobian.
Lemma 16.
Let $(A,B,c)$ be of the specific structure given in (57). Then, for $i=2,\ldots,m$ and $j=2,\ldots,i$
(82) $\frac{\partial Q_{i+2}}{\partial a_{i,j}}=Q_{j}B_{i+2}$
and for $i=3,\ldots,m$ and $j=2,\ldots,i$,
(83) $\left(\frac{\partial}{\partial b_{i,j}}-\frac{\partial}{\partial a_{i,j}}\right)Q_{i+2}=Q_{j}(A_{i+2}-B_{i+2})$
Moreover, for $i=5,\ldots m$,
(84) $\left(\frac{\partial}{\partial b_{i-1,i-1}}-\frac{\partial}{\partial a_{i-1,i-1}}+2\left(\frac{\partial}{\partial b_{i-1,2}}-\frac{\partial}{\partial a_{i-1,2}}\right)+2\frac{\partial}{\partial a_{i,i}}\right)Q_{i+2}=B_{i+2}B_{i-1}Q_{2}+\mathcal{O}(X^{r}),$
where $r={2^{i-1}}$ and also,
(85) $\left(\frac{\partial}{\partial b_{m,m}}-\frac{\partial}{\partial a_{m,m}}+2\left(\frac{\partial}{\partial b_{m,2}}-\frac{\partial}{\partial a_{m,2}}\right)+\frac{\partial}{\partial c_{m+1}}\right)p=Q_{2}(2Q_{2}-B_{m}).$
Proof.
We prove (82) by applying the product rule:
(86) $\frac{\partial Q_{i+2}}{\partial a_{i,j}}=\frac{\partial A_{i+2}}{\partial a_{i,j}}B_{i+2}+A_{i+2}\frac{\partial B_{i+2}}{\partial a_{i,j}}=Q_{j}B_{i+2}.$
Similarly,
(87) $\frac{\partial Q_{i+2}}{\partial b_{i,j}}=Q_{j}A_{i+2}.$
Equation (83) is an immediate consequence of (86) and (87).
For notational convenience, we define the differential operator:
(88) $\mathcal{D}_{k}:=\frac{\partial}{\partial b_{k,k}}-\frac{\partial}{\partial a_{k,k}}+2\left(\frac{\partial}{\partial b_{k,2}}-\frac{\partial}{\partial a_{k,2}}\right),$
for which we derive the auxiliary relation:
(89a) $\mathcal{D}_{k}Q_{k+2}$ $=(A_{k+2}-B_{k+2})(2Q_{2}+Q_{k})$
(89b) $=(Q_{2}-B_{k+1})(Q_{2}+A_{k+1})$
(89c) $=Q_{2}^{2}+Q_{2}(A_{k+1}-B_{k+1})-A_{k+1}B_{k+1}$
(89d) $=2Q_{2}^{2}-Q_{2}B_{k}-Q_{k+1}.$
Here, we have used (83) with $i=k$, $j=2$ and $j=i$, in the first equality, and the fact that for the specific structure given in (57), we have $A_{k+2}-B_{k+2}=Q_{2}-B_{k+1}$.
To prove (85), we use (89) with $k=m$:
(90) $\left(\mathcal{D}_{m}+\frac{\partial}{\partial c_{m+1}}\right)p=2Q_{2}^{2}-Q_{2}B_{m}-Q_{m+1}+Q_{m+1}=2Q_{2}^{2}-Q_{2}B_{m}.$
To prove (84), we use that $\mathcal{D}_{i-1}Q_{2}=\cdots=\mathcal{D}_{i-1}Q_{i}=0$ since these elements are independent of row i1i-1. This, together with the chain rule, yields:
(91a) $\mathcal{D}_{i-1}Q_{i+2}$ $=B_{i+2}\mathcal{D}_{i-1}A_{i+2}+A_{i+2}\mathcal{D}_{i-1}B_{i+2}$
(91b) $=(A_{i+2}+B_{i+2})\mathcal{D}_{i-1}Q_{i+1}$
(91c) $=(A_{i+2}-B_{i+2})\mathcal{D}_{i-1}Q_{i+1}+2B_{i+2}\mathcal{D}_{i-1}Q_{i+1}.$
The final equality is a reformulation which simplifies the following derivation,
(92a) $\left(\mathcal{D}_{i-1}+2\frac{\partial}{\partial a_{i,i}}\right)Q_{i+2}$ $=(A_{i+2}-B_{i+2})\mathcal{D}_{i-1}Q_{i+1}+2B_{i+2}\left(\mathcal{D}_{i-1}Q_{i+1}+Q_{i}\right)$
(92b) $=(A_{i+2}-B_{i+2})(2Q_{2}^{2}-Q_{2}B_{i-1}-Q_{i})+2B_{i+2}\left(2Q_{2}^{2}+Q_{2}B_{i-1}\right),$
where we have used (82) in the first equality. The degree of the first term in the right hand side is given by
(93) $\deg((A_{i+2}-B_{i+2})(2Q_{2}^{2}-Q_{2}B_{i-1}-Q_{i}))$ $=\deg(Q_{i}Q_{i})=2^{i-1},$
and the degree of the second term is given by
(94) $\deg(B_{i+2}B_{i-1}Q_{2})=2^{i-1}+2^{i-4}+1.$
The theorem conclusion (84) follows from Equation (92) and Equation (94).
Proof of Theorem 10
We want to prove the equality in (68); but the upper bound is already given in Corollary 9. To prove that $m^{2}$ is also a lower bound it is sufficient to find one point, i.e., one triplet, with a Jacobian of rank $m^{2}$. This is done by finding particular linear combinations of partial derivatives of the entries in $(A,B,c)$ with $m^{2}$ distinct degrees. We assume the structure given in (57). We construct linear combinations of partial derivatives in 5 different ways. We refer to these linear combinations as Cases 1, 2, 3, 4 and 5.
Case 1: The fact that the evaluation scheme is unreduced means $\deg(Q_{1})=1$, and $\deg(Q_{j})=2^{j-2}$ for $j=2,\ldots,m+2$. This, together with $\deg\left(\frac{\partial p}{\partial c_{j}}\right)=\deg(Q_{j})$, directly implies (58).
For Case 2, we consider the operator
$\mathcal{D}_{i}:=\frac{\partial}{\partial a_{i,j}},$
for $i=2,\ldots,m$ and $j=2,\ldots,i$. Since $a_{i,j}$ does not appear in the first i1i-1 rows, we have $\mathcal{D}_{i}Q_{2}=\cdots=\mathcal{D}_{i}Q_{i+1}=0$. Consequently, the conditions of Lemma 15 are satisfied. Moreover, (82) implies that
(95) $\deg(\mathcal{D}_{i}Q_{i+2})=\deg(Q_{j}B_{i+2})=\deg(Q_{j})+\deg(Q_{i+1})=2^{i-1}+2^{j-2}.$
Combining this with (78) yields
(96a) $\deg(\mathcal{D}_{i}p)$ $=\deg(\mathcal{D}_{i}Q_{m+2})=2^{m-1}+\cdots+2^{i}+\deg(\mathcal{D}_{i}Q_{i+2})$
(96b) $=2^{m-1}+\cdots+2^{i-1}+2^{j-2}.$
We conclude (61).
For Case 3, we consider the operator
$\mathcal{D}_{i}:=\left(\frac{\partial}{\partial b_{i,j}}-\frac{\partial}{\partial b_{i,j}}\right),$
for $i=3,\ldots,m$ and $j=2,\ldots,i-1$. The conditions of Lemma 15 are satisfied by the same reasoning as in Case 2. From (83) we obtain
(97) $\deg(\mathcal{D}_{i}Q_{i+2})=\deg(Q_{j}(A_{i+2}-B_{i+2}))=\deg(-Q_{j}Q_{i})=2^{i-2}+2^{j-2}.$
This, together with (78), implies (64).
For Case 4, we consider the operator
$\mathcal{D}_{m}:=\left(\frac{\partial}{\partial b_{m,m}}-\frac{\partial}{\partial a_{m,m}}+2\left(\frac{\partial}{\partial b_{m,2}}-\frac{\partial}{\partial a_{m,2}}\right)+\frac{\partial}{\partial c_{m+1}}\right).$
From equation (84) we immediately conclude that
(98) $\deg(\mathcal{D}_{m}p)=\deg(\mathcal{D}_{m}Q_{m+2})=\deg(Q_{2}B_{m})=2^{m-3}+1.$
For Case 5, we consider the operator
$\mathcal{D}_{i}:=\left(\frac{\partial}{\partial b_{i-1,i-1}}-\frac{\partial}{\partial a_{i-1,i-1}}+2\left(\frac{\partial}{\partial b_{i-1,2}}-\frac{\partial}{\partial a_{i-1,2}}\right)+2\frac{\partial}{\partial a_{i,i}}\right),$
for $i=5,\ldots,m$. We show that this operator satisfies the assumptions of Lemma 15. Firstly, we observe that $\mathcal{D}_{i}Q_{2}=\ldots=\mathcal{D}_{i}Q_{i}=0$ by the same reasoning as in Case 2. Therefore, it is enough to show that $\deg(\mathcal{D}_{i}Q_{i+1})<\deg(\mathcal{D}_{i}Q_{i+2})$. From (84) and (94) we have that $\deg(\mathcal{D}_{i}Q_{i+2})=2^{i-1}+2^{i-4}+1$. Next, we observe that
(99a) $\deg\left(\mathcal{D}_{i}Q_{i+1}\right)$ $=\deg\left(\left(\frac{\partial}{\partial b_{i-1,i-1}}-\frac{\partial}{\partial a_{i-1,i-1}}+2\left(\frac{\partial}{\partial b_{i-1,2}}-\frac{\partial}{\partial a_{i-1,2}}\right)+2\frac{\partial}{\partial a_{i,i}}\right)Q_{i+1}\right)$
(99b) $=\deg\left((A_{i+1}-B_{i+1})(Q_{2}+Q_{i-1})\right)$
(99c) $=\deg(Q_{i-1}^{2})=2^{i-1}<2^{i-1}+2^{i-4}+1.$
Thus the assumptions of Lemma 15 are satisfied. It follows that
(100a) $\mathcal{D}_{i}p=\mathcal{D}_{i}Q_{i+2}$ $=2^{m-1}+\cdots+2^{i-1}+2^{i-4}+1,\quad i=5,\ldots,m,$
(100b) $=2^{m-1}+\cdots+2^{i}+2^{i-3}+1,\quad i=4,\ldots,m-1.$
To summarize, we have obtained the following degrees through linear combinations of partial derivatives.
  • Case 1:  $0,2^{0},2^{1},\ldots,2^{m}$.
  • Case 2:  $2^{m-1}+\cdots+2^{i}+2^{i-1}+2^{j-2},\;\;\textrm{ for }j=2,\ldots,i,\;\;i=2,\ldots,m$.
  • Case 3:  $2^{m-1}+\cdots+2^{i}+2^{i-2}+2^{j-2},\;\;\;\;\;\textrm{ for }j=2,\ldots,i-1,\;\;i=3,\ldots,m$.
  • Case 4: $2^{m-3}+1$.
  • Case 5: $2^{m-1}+\cdots+2^{i}+2^{i-3}+1,\;\;\;\textrm{ for }i=4,\ldots,m-1.$
These form distinct degrees, in the following way. Firstly, we note that when the degrees are expressed as binary numbers, Case 1 and Case 4 always involve one and two nonzeros respectively, making these degrees distinct. Similarly, when we express the degrees of Case 2, 3 and 5 as binary numbers, they always involve three or more nonzeros, making them distinct from Case 1 and Case 4. For Case 2, Case 3 or Case 5 to coincide, the degrees as binary numbers must have the same number on nonzeros. This corresponds to choosing the same $i$ in each of the formulas, which leads to different degrees. Therefore we conclude that all the above degrees are distinct from each other.
Counting the number of unique degrees, we get: Case 1: $m+1$; Case 2: m(m1)/2m(m-1)/2; Case 3: (m1)(m2)/2(m-1)(m-2)/2; Case 4: $1$; Case 5: m4m-4. In total, we have $m^{2}$ distinct degrees for $m\geq 4$.